# Dual-Objective Reinforcement Learning with Novel Hamilton-Jacobi-Bellman Formulations

### v1

@@ IV-A Augmentations
For the RAA problem, we consider an augmentation of the MDP defined by ¯$\overline{\mathcal{M}}=\langle\overline{\mathcal{{S}}},\mathcal{{A}},\bar{f}\rangle$ with augmented states $\overline{\mathcal{{S}}}=\mathcal{{S}}\times\mathcal{{Y}}\times\mathcal{{Z}}$, where $\mathcal{{Y}}:=\{r(s)\mid s\in\mathcal{{S}}\}$ and $\mathcal{{Z}}:=\{q(s)\mid s\in\mathcal{{S}}\}$. Note the actions $\mathcal{{A}}$ remain unchanged. For any initial state $s$, let the augmented states be initialized as $y=r(s)$ and $z=q(s)$, and let the transition $\bar{f}$ of $\overline{\mathcal{M}}$ be given by
$s_{t+1}^{\bar{\pi}}$ $=f\left(s_{t}^{\bar{\pi}},\bar{\pi}\left(s_{t}^{\bar{\pi}},y_{t}^{\bar{\pi}},z_{t}^{\bar{\pi}}\right)\right),$
$y_{t+1}^{\bar{\pi}}$ $=\max\left\{r\left(s_{t+1}^{\bar{\pi}}\right),y_{t}^{\bar{\pi}}\right\},$
$z_{t+1}^{\bar{\pi}}$ $=\min\left\{q\left(s_{t+1}^{\bar{\pi}}\right),z_{t}^{\bar{\pi}}\right\},$
@@
so that $y_{t}$ and $z_{t}$ track the best reward and worst penalty yet seen. Hence, a policy $\bar{\pi}:\overline{\mathcal{{S}}}\to\mathcal{{A}}$ for $\overline{\mathcal{M}}$ may leverage information regarding the history of the trajectory.
@@
For the RR problem, we augment the system similarly, except $\mathcal{{Y}}:=\{r_{1}(s)\mid s\in\mathcal{{S}}\}$, $\mathcal{{Z}}:=\{r_{2}(s)\mid s\in\mathcal{{S}}\}$, and
$s_{t+1}^{\bar{\pi}}$ $=f\left(s_{t}^{\bar{\pi}},\bar{\pi}\left(s_{t}^{\bar{\pi}},y_{t}^{\bar{\pi}},z_{t}^{\bar{\pi}}\right)\right),$
$y_{t+1}^{\bar{\pi}}$ $=\max\left\{r_{1}\left(s_{t+1}^{\bar{\pi}}\right),y_{t}^{\bar{\pi}}\right\},$
$z_{t+1}^{\bar{\pi}}$ $=\max\left\{r_{2}\left(s_{t+1}^{\bar{\pi}}\right),z_{t}^{\bar{\pi}}\right\}.$
V Optimal Policies for RAA and RR by Value Decomposition
@@
We now discuss our first theoretical contributions.
V-A Decomposition of RAA into avoid and reach-avoid problems
@@
Our main theoretical result for the RAA problem shows that we can solve this problem by first solving the avoid problem corresponding to the negated penalty $q(s)$ to obtain the optimal value function $V_{\textup{A}}^{*}(s)$ and then solving a reach-avoid problem with the negated penalty $q(s)$ and a modified reward $r_{\textup{RAA}}(s)$.
@@
Proof.
See section A of the Supplementary Material. ∎
@@
V-B Decomposition of the RR problem into three reach problems
@@
$\max_{\bar{\pi}}V_{\textup{RAA}}^{\bar{\pi}}(s)=\max_{\pi}\max_{t}r_{\textup{RR}}\left(s_{t}^{\pi}\right),$ (2)
where
$r_{\textup{RR}}(s):=\max\left\{\min\left\{r_{1}(s),V_{\textup{R2}}^{*}(s)\right\},\min\left\{r_{2}(s),V_{\textup{R1}}^{*}(s)\right\}\right\},$
with
$V_{\textup{R1}}^{*}(s):=\max_{\pi}\max_{t}r_{1}\left(s_{t}^{\pi}\right)\text{ and }V_{\textup{R2}}^{*}(s):=\max_{\pi}\max_{t}r_{2}\left(s_{t}^{\pi}\right).$
Proof.
See section B of the Supplementary Material. ∎
@@
V-C Optimality of the augmented problems
The following theoretical result demonstrates that the augmentation is optimal for the original problem in that no further historical information can be used to improve performance.
@@ V-C Optimality of the augmented problems
$\max_{\bar{\pi}}V_{\textup{RAA}}^{\bar{\pi}}(s)$ $=\max_{a_{0},a_{1},\dots}\min\left\{\max_{t}r(s_{t}),\min_{t}q(s_{t})\right\},$
@@
$\max_{\bar{\pi}}V_{\textup{RR}}^{\bar{\pi}}(s)$ $=\max_{a_{0},a_{1},\dots}\min\left\{\max_{t}r_{1}(s_{t}),\max_{t}r_{2}(s_{t})\right\}$
@@
VI DO-HJ-PPO: Solving RAA and RR with RL
@@
In the previous sections, we demonstrated that the RAA and RR problems can be solved through decomposition of the values into reach, avoid, and reach-avoid problems. Because each of these problems has a corresponding Bellman relation (see [18, 19, 21]), these decompositions allow us to solve the dual-objective problems by minimally modifying existing RL methods. We make a few assumptions in the derivation that would limit performance and generalization, namely, the determinism of the values as well as access to the decomposed values (by solving them beforehand). In this section, we propose relaxations to the RR and RAA theory and devise a custom variant of Proximal Policy Optimization, DO-HJ-PPO, to solve this broader class of problems, and demonstrate its performance.
VI-A Stochastic Reach-Avoid Bellman Equation
@@
It is well known that the most performative RL methods allow for stochastic learning. In [21], the Stochastic Reachability Bellman Equation (SRBE) is described for Reach problems and used to design a specialized PPO algorithm. In this section we proceed by closely following this work, modifying the SRBE into a Stochastic Reach-Avoid Bellman Equation (SRABE). Using Theorems 1 and 2, the SRBE and SRABE offer the necessary tools for designing a PPO variant for solving the RR and RAA problems.
We define $\tilde{V}_{\textup{RAA}}^{\pi}$ to be the solution to the following Bellman equation (SRABE):
$\tilde{V}_{\textup{RAA}}^{\pi}$ $(s)=$
$\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\right]$
The corresponding action-value function is
We then have the following proposition:
Proposition 1.
For each $s\in\mathcal{{S}}$ and every $\theta\in\mathbb{R}^{n_{p}}$, we have
with the bracketed term equal to $1$ if the proposition inside is true and 00 otherwise.
Figure 3: Algorithm Comparisons for RR and RAA tasks. We evaluate our method and relevant baselines on 1,000 trajectories for both the Reach-Reach (RR) and Reach-Always-Avoid (RAA) problems for Hopper and F16 environments. In the RR tasks, we compare against a decomposed version of the problem (DSTL) and several variants of CPPO. Our method consistently reaches both target regions with a higher success rate and fewer steps on average. Notably, CPPOv1 and CPPOv2 fail to achieve any successful trajectories in the RR task, whereas CPPOv3 shows improved—but still limited—performance. For the RAA tasks, we compare our approach against Constrained PPO (CPPO) and standard reach-avoid baselines. Our method achieves a higher success rate while requiring a lower average number of steps to reach success.
Following [19], we then define the discounted value and action-value functions with $\gamma\in[0,1)$:
$\tilde{V}_{\textup{RAA}}^{\gamma,\pi}(s)=(1-\gamma)\min\left\{r_{\textup{RAA}}(s),q(s)\right\}$
$\qquad+\gamma\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\gamma,\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\right],$
$\tilde{Q}_{\textup{RAA}}^{\gamma,\pi}(s,a)=(1-\gamma)\min\left\{r_{\textup{RAA}}(s),q(s)\right\}$
$\qquad+\gamma\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\gamma,\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}.$
The PPO advantage function is then given by $\tilde{A}_{\textup{RAA}}^{\pi}=\tilde{Q}_{\textup{RAA}}-\tilde{V}_{\textup{RAA}}$ [23].
VI-B Algorithm
We introduce DO-HJ-PPO in Algorithm 1, a unified PPO-based algorithm for solving the Reach-Always-Avoid (RAA) and Reach-Reach (RR) problems, which builds on the SRABE and SRBE formulations with minimal modifications to the standard PPO framework. Note that in the case of a reset while learning, augmented state variables are set to the initial values described in section IV.
Algorithm 1 : DO-HJ-PPO (Actor-Critic)
1:Define Composed Actor and Critic $\tilde{Q}$
2:Define Decomposed Actor(s) and Critic(s) $\tilde{Q}_{i}$
3:for $k=0,1,\ldots$ do
4:     for $t=0$ to $T-1$ do
5:         Sample trajectories for $\tau_{t}:\{\hat{s}_{t},a_{t},\hat{s}_{t+1}\}$
6:         Define $\tilde{r}(s_{t})$ with Decomposed Critics $\tilde{Q}_{i}(s_{t})$ (Theorems  1 & 2)
7:         Composed Critic update:
$\omega\leftarrow\omega-\beta_{k}\nabla_{\omega}\tilde{Q}(\tau_{t})\cdot\left(\tilde{Q}(\tau_{t})-B^{\gamma}[\tilde{Q};\tilde{r}](\tau_{t})\right)$
8:         Compute Bellman-GAE $A_{HJ}^{\lambda}$ with $B^{\gamma}$
9:         (Standard) update Composed Actor
10:         Decomposed Critic update(s):
11:         Compute Bellman-GAE $A_{i}^{\lambda}$ with $B^{\gamma}_{i}$
12:         (Standard) update Decomposed Actor(s)
13:     end for
14:end for
15:return parameter $\theta$, $\omega$
In Algorithm 1, $B^{\gamma}$ and $B^{\gamma}_{i}$ represent the composed and decomposed Bellman updates for the users choice of problem (RAA or RR). The Bellman update $B^{\gamma}[\tilde{Q};\tilde{r}]$ differs for the RAA task and RR task, depending on the corresponding the reach, avoid, and reach-avoid decomposition. Beyond the modified Bellman updates, the learning proceeds in the nominal PPO fashion, in which the actively-learned decomposed values are used to define the composed value (see Supplementary).
VII Experiments
@@
We first demonstrate the theoretical results (Theorems 1 and 2) through a simple 2D grid-world experiment using Double Deep Q-Networks (DDQN) (Figure 1). Additional experimental details are provided in the Supplementary Material. On the left, we compare the optimal value functions learned under the classic RA formulation with those from the RAA setting. In the RA scenario, trajectories successfully avoid the obstacle but may terminate in regions from which future collisions are inevitable. On the right, we consider a similar environment but with two reward targets. Here, the RR formulation induces trajectories that visit both targets, unlike simple reach tasks in which the agent halts. These qualitative results highlight the behavioral distinctions induced by the RAA and RR objectives compared to their simpler counterparts.
To evaluate the method under more complex and less structured conditions, we extend our analysis to continuous control settings using our algorithm DO-HJ-PPO. Specifically, we apply DO-HJ-PPO to RAA and RR tasks in the Hopper and F16 environments. For the Hopper, two high targets and floor and wall obstacles are defined with respect to its head, and in the F16, the targets are defined by regions to fly through, while obstacles are defined by geofences which create a boxed flight corridor. We compare against both STL (DSTL) and contrained PPO (CPPO) baselines (see supplementary material). Empirically, DO-HJ-PPO performs equivalently at worst and more often at a significantly higher ability, scored in metrics of task success percentage and steps to achieve the task, indicating that DO-HJ-PPO more reliably and rapidly solves the given tasks. These results underscore the challenging nature of composing multiple objectives using traditional baselines while in contrast, our method provides a more robust and direct solution for handling such complex compositional tasks, with less required tuning.
VIII Conclusion
In this brief, we introduced two novel Bellman formulations for new problems (RAA and RR) which generalize those in recent publications. We prove decomposition results for these problems that allow us to break them into simpler Bellman problems, which can then be composed to obtain the value functions and corresponding optimal policies. We use these results to design a PPO-based algorithm for practical solution of RAA and RR. More broadly, this work provides a road-map to extend the range of Bellman formulations that can be solved, via decomposing higher-level problems into lower-level ones. By solving the RAA and RR, we add two new ingredients to the list of solvable problems that can be leveraged toward this end.
Supplementary Material

### v2

@@ 5.1 Augmentation of the RAA Problem
We consider an augmentation of the MDP defined by $\overline{\mathcal{M}}=\langle\overline{\mathcal{{S}}},\mathcal{{A}},f\rangle$ consisting of augmented states $\overline{\mathcal{{S}}}=\mathcal{{S}}\times\mathcal{{Y}}\times\mathcal{{Z}}$ and the same actions $\mathcal{{A}}$. For any initial state $s$, let the augmented states be initialized as $y=r(s)$ and $z=q(s)$, and let the transition of $\overline{\mathcal{M}}$ be defined by
$s_{t+1}^{\bar{\pi}}=f\left(s_{t}^{\bar{\pi}},\bar{\pi}\left(s_{t}^{\bar{\pi}},y_{t}^{\bar{\pi}},z_{t}^{\bar{\pi}}\right)\right);\quad y_{t+1}^{\bar{\pi}}=\max\left\{r\left(s_{t+1}^{\bar{\pi}}\right),y_{t}^{\bar{\pi}}\right\};\quad z_{t+1}^{\bar{\pi}}=\min\left\{q\left(s_{t+1}^{\bar{\pi}}\right),z_{t}^{\bar{\pi}}\right\},$
@@
such that $y_{t}$ and $z_{t}$ track the best reward and worst penalty up to any point. Hence, the policy for $\overline{\mathcal{M}}$ given by $\bar{\pi}:\overline{\mathcal{{S}}}\to\mathcal{{A}}$ may now consider information regarding the history of the trajectory.
By definition, the RAA value for $\overline{\mathcal{M}}$,
$V_{\textup{RAA}}^{\bar{\pi}}(s)=\min\left\{\max_{t}r(s_{t}^{\bar{\pi}}),\min_{t}q(s_{t}^{\bar{\pi}})\right\},$
is equivalent to that of $\mathcal{M}$ except that it allows for a policy $\bar{\pi}$ which has access to historical information. We seek to find $\bar{\pi}$ that maximizes this value.
5.2 Augmentation of the RR Problem
@@
For the Reach-Reach problem, we augment the system similarly, except that $z_{t}$ is updated using a $\max$ operation instead of a $\min$:
$s_{t+1}^{\bar{\pi}}=f\left(s_{t}^{\bar{\pi}},\bar{\pi}\left(s_{t}^{\bar{\pi}},y_{t}^{\bar{\pi}},z_{t}^{\bar{\pi}}\right)\right);\quad y_{t+1}^{\bar{\pi}}=\max\left\{r_{1}\left(s_{t+1}^{\bar{\pi}}\right),y_{t}^{\bar{\pi}}\right\};\quad z_{t+1}^{\bar{\pi}}=\max\left\{r_{2}\left(s_{t+1}^{\bar{\pi}}\right),z_{t}^{\bar{\pi}}\right\}.$
Again, by definition,
$V_{\textup{RR}}^{\bar{\pi}}(s)=\min\left\{\max_{t}r_{1}(s_{t}^{\bar{\pi}}),\max_{t}r_{2}(s_{t}^{\bar{\pi}})\right\}.$
The RR problem is again to find an augmented policy $\bar{\pi}$ which maximizes this value.
6 Optimal Policies for RAA and RR by Value Decomposition
@@
We now discuss our first theoretical contributions. We refer the reader to the appendix for the proofs of the theorems.
6.1 Decomposition of RAA into avoid and reach-avoid problems
@@
Our main theoretical result for the RAA problem shows that we can solve this problem by first solving the avoid problem corresponding to the penalty $q(s)$ to obtain the optimal value function $V_{\textup{A}}^{*}(s)$ and then solving a reach-avoid problem with the negated penalty function $q(s)$ and a modified reward function $r_{\textup{RAA}}(s)$.
@@
This decomposition is significant, as methods customized to solving avoid and reach-avoid problems were recently explored in Fisac et al. (2019); Hsu et al. (2021); So et al. (2024); So and Fan (2023), allowing us to effectively solve the optimization problem defining $V_{\textup{A}}^{*}(s)$ as well as the optimization problem that defines the right-hand-side of 1.
@@
Readers familiar with temporal logic (TL) may be interested in how these decompositions relate to decompositions of predicates in TL. We discuss the distinction between these two classes of decompositions in Sec. L of the Appendix, and how the TL predicate algebra is insufficient for safe optimal control.
6.2 Decomposition of the RR problem into three reach problems
@@
$\max_{\bar{\pi}}V_{\textup{RR}}^{\bar{\pi}}(s)=\max_{\pi}\max_{t}r_{\textup{RR}}\left(s_{t}^{\pi}\right),$ (2)
where $r_{\textup{RR}}(s):=\max\left\{\min\left\{r_{1}(s),V_{\textup{R2}}^{*}(s)\right\},\min\left\{r_{2}(s),V_{\textup{R1}}^{*}(s)\right\}\right\}$, with
$V_{\textup{R1}}^{*}(s):=\max_{\pi}\max_{t}r_{1}\left(s_{t}^{\pi}\right),\quad V_{\textup{R2}}^{*}(s):=\max_{\pi}\max_{t}r_{2}\left(s_{t}^{\pi}\right).$
@@
6.3 Optimality of the augmented problems
We previously motivated the choice to consider an augmented MDP $\overline{\mathcal{M}}$ over the original MDP in the context of the RAA and RR problems. In this section, we justify our particular choice of augmentation. Indeed, the following theoretical result shows that further augmenting the states with additional historical information cannot improve performance under the optimal policy.
@@ 6.3 Optimality of the augmented problems
$\max_{\pi}V_{\textup{RAA}}^{\pi}(s)\leq\max_{\bar{\pi}}V_{\textup{RAA}}^{\bar{\pi}}(s)=\max_{a_{0},a_{1},\dots}\min\left\{\max_{t}r(s_{t}),\min_{t}q(s_{t})\right\},$
@@
$\max_{\pi}V_{\textup{RR}}^{\pi}(s)\leq\max_{\bar{\pi}}V_{\textup{RR}}^{\bar{\pi}}(s)=\max_{a_{0},a_{1},\dots}\min\left\{\max_{t}r_{1}(s_{t}),\max_{t}r_{2}(s_{t})\right\}$
@@
The terms on the right of the lines above reflect the best possible sequence of actions to solve the RAA or RR problem, and the theorem states that the optimal augmented policy achieves that value, represented by the middle terms. This value will generally be less than or equal to the outcome from using a non-augmented policy, represented by the terms on the left.
7 DOHJ-PPO: Solving RAA and RR with RL
@@
In the previous sections, we demonstrated that the RAA and RR problems can be solved through decomposition of the values into formulations amenable to existing RL methods. However, we make a few assumptions in the derivation that would limit performance and generalization, namely, the determinism of the values as well as access to the decomposed values (by solving them beforehand). In this section, we propose relaxations to the RR and RAA theory and devise a custom variant of Proximal Policy Optimization, DOHJ-PPO, to solve this broader class of problems, and demonstrate its performance.
7.1 Stochastic Reach-Avoid Bellman Equation
@@
It is well known that the most performative RL methods allow for stochastic learning. In So et al. (2024), the Stochastic Reachability Bellman Equation (SRBE) is described for Reach problems and used to design a specialized PPO algorithm. We first generalize this notion to a Stochastic Reach-Avoid Bellman Equation (SRABE). Using Theorems 1 and 2, the SRBE and SRABE offer the necessary tools for designing a PPO variant for solving the RR and RAA problems.
By anology to the SRBE, the SRABE is given by
$\tilde{V}_{\textup{RAA}}^{\pi}(s)=\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\right].$ (SRABE)
More rigorously, we actually consider the discounted SRABE, which is contractive, and the corresponding quality function below in the limit $\gamma\to 1^{-}$ (as in Hsu et al. (2021)),
$\tilde{V}_{\textup{RAA}}^{\gamma,\pi}(s)=(1-\gamma)\min\left\{r_{\textup{RAA}}(s),q(s)\right\}+\gamma\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\gamma,\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\right].$
$\tilde{Q}_{\textup{RAA}}^{\gamma,\pi}(s,a)=(1-\gamma)\min\left\{r_{\textup{RAA}}(s),q(s)\right\}+\gamma\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\gamma,\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}.$
Theoretically speaking, the use of the SRABE is justified by Theorem 4 in the Appendix. With this action-value function we then follow So et al. to derive the corresponding policy gradient result with an augmented version of the dynamics; for details, see Prop. X in the Appendix. The PPO advantage function is then given by $\tilde{A}_{\textup{RAA}}^{\pi}=\tilde{Q}_{\textup{RAA}}-\tilde{V}_{\textup{RAA}}$ Schulman et al. (2017). Although, this approximation may be poor in highly noisy settings, we show this approach yields conservative estimates of the value with stochastic policies (Appendix sec. D), and validate it empirically with stochastic dynamics in Sec. 8.3.
7.2 Algorithm
We introduce DOHJ-PPO for solving the RAA and RR problems, which integrates the SRABE and SRBE via three minimal modifications to PPO Schulman et al. (2017) (see appendix for more).
Additional actor and critics are introduced to represent the decomposed objectives. Per Theorems 1 and 2, one may know that the RAA and RR values are given by a composition of the simpler R, A and RA values. Therefore, we learn these decompositions with their own networks and integrate them into the composed actor and critic training, namely via the GAE and target with the special RAA and RR reward functions in Theorems 1 and 2.
The composed actor and critic are learned concurrently to the decomposed actor and critics by bootstrapping the current values Rather than learning the decomposed and composed representations sequentially, DOHJ-PPO bootstraps to learn them simultaneously. Namely, at each iteration, we rollout trajectories for composed and decomposed updates with each actor. In the update of the composed representation specifically, the decomposed values are inferred from the current decomposed critic(s) along the composed trajectories. This design choice allows us to couple the on-policy learning of PPO in the following way.
Trajectories for training the decomposed actor and critic(s) are initialized with states sampled from the composed trajectories, which we refer to as coupled resets. While it is possible to estimate the decomposed objectives independently—i.e., prior to solving the composed task—this approach might lead to inaccurate or irrelevant value estimates in on-policy settings. For example, in the RAA problem, the avoid decomposition will solely prioritize avoiding penalties and, hence, might converge to an optimal strategy within a reward-irrelevant region, misaligned with the overall task.
8 Experiments
8.1 DQN Demonstration
@@
We begin by demonstrating the utility of our theoretical results (Theorems 1 and 2) through a simple 2D grid-world experiment using DQN (Figure 2). In this environment, the agent can move left, right, or remain stationary, while drifting upward at a constant rate. Throughout, reward regions are shown in blue and penalty regions in red. On the left, we compare the optimal value functions learned under the classic Reach-Avoid (RA) formulation with those from the Reach-Always-Avoid (RAA) setting. In the RA scenario, trajectories successfully avoid the obstacle but may terminate in regions from which future collisions are inevitable, as there is no incentive to consider what happens after reaching the minimum reward threshold. In contrast, under the RAA formulation, where the objective involves maximizing cumulative reward while accounting for future penalties (as per Theorem 1), the agent learns to reach the target while remaining in safe regions thereafter. On the right, we consider a similar environment without obstacles but with two distinct targets. Here, the Reach-Reach (RR) formulation induces trajectories that visit both targets, unlike simple reach tasks in which the agent halts after reaching a single goal. These qualitative results highlight the behavioral distinctions induced by the RAA and RR objectives compared to their simpler counterparts. Additional algorithmic and experimental details are provided in the Apendix.
Figure 4: Success ($\rightarrow$) and Partial Success ($\rightarrow$) in RAA and RR Tasks for DOHJ-PPO and Baselines. We evaluate DOHJ-PPO in black against baselines over 1,000 trajectories in the Hopper, F16, SafetyGym and HalfCheetah environments. In the first and third row, the Partial Success percentage of each algorithm is given, defined by the number of trajectories to achieve one objective (reaching or always-avoiding in the RAA, reaching either in the $R$). In the second and fourth rows, SUCCESS percentage is given, defined by the number of trajectories to achieve both objectives. Most baselines achieve partial success, however, few achieve total success as the environment becomes more difficult, underscoring the difficulty of balancing objectives in RL.
8.2 Continuous Control Tasks with DOHJ-PPO
To evaluate the method under more complex and less structured conditions, we extend our analysis to continuous control settings. Specifically, we consider RAA and RR tasks in the Hopper, F16, SafetyGym, and HalfCheetah environments, depicted in Figure 1. In the RAA tasks, the penalty function generally characterizes regions of states where the agent (or its body parts) is intended to avoid, while the reward characterizes regions of states where the agent is intended to reach.
As baselines, we compare DOHJ-PPO against a variety of classes of RL algorithms. We include several augmented Lagrangian methods which transform constraints (either for reaching both or always avoiding) into mixed objectives, such as Constrained PPO (CPPO) Achiam et al. (2017b), PPO-LAG Ray et al. (2019), P2BPO Dey et al. (2024), and LOGBAR Zhang et al. (2024). Additionally, we include three HJ-RL baselines designed for the previous R and RA problems, RESPO Ganai et al. (2023), RCPPO So et al. (2024) and RA Hsu et al. (2021). Lastly, we also include a few methods based on approaches in STL/LTL-RL and MORL, including a decomposed STL (D-STL) PPO, a sparse-reward STL PPO (SPARSE) and a MORL-based PPO. All algorithms are trained on random initial conditions and then evaluated on new random initial conditions within distribution. To quantify performance of the dual-objective tasks, we measure (1) the percent of trajectories which achieve at least both tasks successfully, (2) the percent of trajectories which achieve at least one of the tasks (dubbed partial success), and (3) the mean steps in each trajectory until success.
Empirically, we find that our method performs at the top-level, achieving first or second place among all tasks and environments (Figure 4). In fact, as the multi-target (RR) or safe-achievement (RAA) tasks become more complex (e.g. the HalfCheetah), our algorithm increasingly dominates the 10 state-of-the-art baselines. Note, that almost all algorithms can achieve partial success at a high rate in each dual-objective task, highlighting the difficulty of mixed or competing objectives, particularly with discounted-sum rewards. Moreover, DOHJ-PPO is the sole performant algorithm in both RAA and RR tasks, displaying the fastest achievement times across tasks (see appendix).
These results underscore the challenging nature of composing multiple satisfaction objectives using traditional baselines with discounted-sum rewards. In contrast, DOHJ-PPO provides a direct and robust solution to handling these complex tasks, with little to no tuning. Our algorithm enjoys these benefits because of the structure of the novel Bellman updates, which propagate the extreme (maximum and minimum) values as opposed to the short-term average (discounted-sum) values.
Figure 5: Success ($\uparrow$) in the HalfCheetah RAA and RR Tasks with Increasingly Stochastic Dynamics. We plot the learning trajectories of DOHJ-PPO in black and the top baselines for the HalfCheetah environment with an affine Gaussian noise added to the dynamics. Task achievement (success) is given by the percentage of 256 trajectories that either reach the target and always-avoid the obstacles or reach both targets (corresponding to $V_{\mathrm{RAA}}>0$ and $V_{\mathrm{RR}}>0$). Each column corresponds to a different scale of noise – null, low (0.5), moderate (1.) and high (2.) – which is added to the velocities and angular velocities of the HalfCheetah dynamics. In the RAA task, DOHJ-PPO outperforms all baselines up to the highest noise settings where all algorithms perform equivalently poorly. In the RR task, DOHJ-PPO outperforms all algorithms significantly. In summary, this ablation demonstrates the robustness of DOHJ-PPO to certain stochasticity in the dynamics and the validity of the SRBE and SRABE approximations.
8.3 Comparison in Stochastic Dynamics
To design an algorithm robust to randomness, DOHJ-PPO employs the SRBE and SRABE discussed in Sec. 7.1 in place of their analogous deterministic forms. This choice equates to an approximation of the decompositional results (Thms. 1 and  2) that interchanges the extrema and expectation operators. The empirical results in Fig. 4 justify this approximation with stochastic policies, however, this noise is introduced for exploration and ultimately attenuated in training. To interrogate the behavior of DOHJ-PPO with stochastic dynamics, we inject affine Gaussian noise into the evolution of the HalfCheetah dynamics for both RAA and RR tasks. Note, only the velocities and angular velocities of the agent are perturbed to protect contact physics. We compare our algorithm against the top three baselines in each task along a scale of standard deviations of low (0.5), moderate (1.) and high (2.) quantity, plotting the maximum learning curve over three seeds in Fig. 5.
In this ablation, we find that the proposed approach, using the novel Bellman equations with stochastic relaxations, offers a significant performance improvement even in the face of significant noise. In the RAA task, DOHJ-PPO dominates the top performing baselines with a 8%-22% peak-performance gap between it and the second best algorithm (and is the fastest to peak-performance) for moderate noise levels, beyond which all algorithms perform equally poorly. In the RR case, we find an even starker result, with all but one experiment demonstrating a >30% improvement in peak-performance even in the high-noise regimes, with an exception of the moderate noise case where DOHJ-PPO still performs >15% than the best baseline, DSTL. Interestingly, DOHJ-PPO is the slower than DSTL to peak performance, but performs twice as well at best in three of the four settings. These results demonstrate that despite certain highly-noisy dynamics DOHJ-PPO is competitive at worst and optimal in majority. See Appendix Sec. D for further analysis and discussion of the usage of the SRBE and SRABE approximations.
9 Conclusions
In this work, we introduced two novel Bellman formulations for new problems (RAA and RR) which generalize those considered in several recent publications. We derive decomposition results to break them into simpler Bellman equations, which can then be composed to obtain the corresponding value functions and optimal policies. We use these results to design DOHJ-PPO, which shows to be the most performant and balanced algorithm in safe-arrival and multi-target achievement. DOHJ-PPO employs the stochastic relaxations of the simpler Bellman equations (the SRBE and SRABE), for which we offer rigorous justification and empirical validation in the case of stochastic policies. As expectation and extrema operations do not commute, more work is needed to provide guarantees under stochastic dynamics. Nonetheless, we demonstrate through an artificial ablation that DOHJ-PPO can be successful in the face of certain dynamic randomness. With regard to more complex objectives, it appears one might employ our results to iteratively decompose layered objectives corresponding to temporal logic specifications into a graph of Bellman values. However, doing so would require deriving generalized decomposition principles for nontrivial compositions of logical operations. Moreover, a practical algorithm for solving the decomposed graph of values might benefit from a more efficient representation, mechanisms to guarantee convergence, and heuristics to improve sampling efficiency, but we leave this to future work. By solving the RAA and RR values, this work provides a road-map to extend complex Bellman formulations, via decomposing higher-level problems into lower-level ones, establishing a foundation for nuanced tasks in real-world environments and safe RL.
Appendix